-- CarmeloArcidiacono - 08 Oct 2011

Arrival at LBT at ~11.00

Carmelo, Lorenzo, Alfio

Sunny, temperature ~ 0 deg Celsius

OPERATIONS

DAYTIME 12.15

FLAO W1 & adcecdx startup, everything ok

SOFTWARE BUGS.

1 W Software without adsecdx running

2 Controller CCD 47

Board setup PISCES RETRO

3. BUG CAMERA LENS IN COLD BCU (3deg C)

FLAO2 start

adcsecsx ok

W2 : X stage started up only at the third attempt (homing failing)


TEST on FLAO1

Centering camera lens

Centering stages

Filter wheel

47

39

rotator

ADC

EVERYTHING ok


Looking for targets

http://catserver.ing.iac.es/staralt/

Riddle 15 19 11 09 +14 50 04 MAG 12,13

Berkeley 82 19 11 20 +13 07 06 MAG 10.4 13.0

Kronberger 68 20 00 36 +30 35 23 MAG 9.2 13.00

Check for the web cam. We did not found it.

Check for if USB port available on PISCES filter motor: we didn't find available ports.

Moreover the optical encoder was removed. And therefore USB and WebCam are useles.

Temperature

FLAO1 (DX) Electronic Boxes

Initial Temperature at 12.00 5deg (ambient 5deg)

Final Temperature at 18.00 15deg (ambient 5deg)

NIGHT

AO814 R_Mag = 7.85 ELEVATION 72

CCD47 Images

Seeing measurements on the CCD47

20111009_025516 @850 ì40@ (600-1000 in front

20111009_025716

camera lens temperature 6.4deg!!!!

TEST AUTOGAIN

3 groups BIN 1 400 modes 030739

3.1, 1.7, 1.1

OPTICAL LOOP DATA FOR TEST 20111009_030739

2.5, 2.1, 1.2

We try switching off the mirror ventilation

(Seeing is ~1.3)

Close loop Pause doesn't work (was intentionally removed!!!)

CCD47 @BIN1

20111009_031730 CCD47 10Frame

But it gave errors WFS arbitrator

Fatto a manina salvato in psf.fits della cartella 20111009_031730

TEST OF PRESET USING IRAF

SCRIPT IDL FOR DOUG's observe IDL routine

# Test of observing script

# 08Oct2011 CA

preset, OBJECT=AO814, MODE=ace, GS=1, RS=0, PA=90

auto_collimate

runao

Initially we had a problem because of the bpp12 version of the LBT software

variables not properly configured, problem solved

and script properly working.

At the end of the script the AO was closed and stable.

We checked also the offset command that was successful

moving the stages of two arcsec pausing the loop and restarting.

Resume of acquisitions

We cannot use PISCES because it is not cold. We use the TV.

The pupil mismatch comment is due to the digital shift of the pupils operated in September. elab_lib should check pupfrom to understand what the original pupils are.

The autogain are extremely high. Why?

The FWHM of the PSF on the 47 is 3px=40mas, actually slightly elongated

TrackNo RefStar Mag El Wind DIMM/OL Rec bin #mod freq gain mod nph AntiDrift SR disturb SN notes
20111009_025516 AO814 8.9 72 9 1.65 -1.00   1 -1 197 -1.0 -1.0 -1.0 0 851 ON -1.0 ONSKY NO - Pupils not on target
20111009_025716 AO814 9.0 72 5 1.71 -1.00   1 -1 197 -1.0 -1.0 -1.0 0 748 ON -1.0 ONSKY NO - Pupils not on target
20111009_030739 AO814 9.0 74 5 NaN 1.03 092604 1 400 990 3.1 1.7 1.1 3 148 ON -1.0 ONSKY NO - Pupils mismatch
20111009_031730 AO814 8.9 75 4 NaN 0.94 092604 1 400 990 2.1 1.5 1.2 3 156 ON -1.0 ONSKY NO - Pupils mismatch
20111009_031759 AO814 9.0 75 6 NaN 1.21 092604 1 400 990 2.1 1.5 1.2 3 146 ON -1.0 ONSKY NO - Pupils mismatch
20111009_032053 AO814 8.9 76 6 NaN 0.90 092604 1 400 990 2.1 1.5 1.2 3 158 ON -1.0 ONSKY NO - Pupils mismatch
Topic revision: r6 - 11 Sep 2012, EnricoPinna
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback